Assume that the probability of a male birth is 1/2. Find the probability that in a family of 4 children there will be:


a) at least 1 boy and at least 1 girl. 


1
Expert's answer
2022-01-09T09:51:43-0500

"P(0\; boy)=0.5^4=0.0625."

"P(0\;girl)=0.5^4=0.0625."

"P( at\;least\;1\;boy\;and\;at\;least\;1\;girl)=1-P(0\;boy)-P(0\;girl)="

"=1-0.0625-0.0625=0.875."